OpenStudy (anonymous):

Two lightweight metal spheres are suspended near each other from insulating threads. One sphere has a net charge; the other sphere has no net charge. The spheres will: A. Attract each other B. Exert no electrostatic force on each other C. Repel each other D. Do any of these things depending on the sign of the charge on the one sphere

OpenStudy (anonymous):

let's do elctrostatic force \[E_0 \frac{Q_0 Q_1}{r^2}\] Q_0 and Q_1 are charge on each sphere let one of them be no charge Q_0=0 \[E_0 \frac{0* Q_1}{r^2}=0\]

OpenStudy (anonymous):

so does that mean B. exert no electrostatic force on each other?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

awesome...can u look at my other post?

OpenStudy (anonymous):

hey...one of the answers of the 2 questions is wrong

OpenStudy (anonymous):

any ideas?

OpenStudy (anonymous):

can't be this one

OpenStudy (jamesj):

Actually, the answer proposed for this one is wrong. Do watch this http://ocw.mit.edu/courses/physics/8-02-electricity-and-magnetism-spring-2002/video-lectures/lecture-1-what-holds-our-world-together/ from minute 7. You see that around minute 9 the non-charged balloon does move towards the charged rod. Something similar is happening with the spheres. Suppose the sphere is positively charged. It doesn't make a big difference to what follows; just swap negative and positive if the sphere is negatively charged. The positively charged sphere causes the other sphere to be polarized, with negative charge on the side closest to the first sphere. Now the second sphere moves towards the first sphere.

OpenStudy (anonymous):

so does that mean the question with the negatively charged rod was correct with positive?

OpenStudy (jamesj):

Oh yes, the answer to that other question doesn't change. Exactly the same principle is at work.

OpenStudy (anonymous):

well we did change that answer from positive to cannot be determined is why i ask

OpenStudy (jamesj):

did we ... let me look at it again.

OpenStudy (anonymous):

oh, just realized the formula I used only works in electrostatic situation. Since it is metal sphere(conductor) we can't use it.

OpenStudy (anonymous):

the answer is A?